The mean value of land and buildings per acre from a sample of farms is $1600, with a standard deviation of $300. The data set has a bell-shaped distribution. Assume the number of farms in the sample is 78.

a) Estimate the number of farms whose land and building values per acre are between $1300 and $1900. 

 

 

b) What percent of the farms have a value below $1900?

 




1
Expert's answer
2021-02-02T05:04:39-0500

a) "P(1300<X<1900)=P(\frac{1300-1600}{300}<Z<\frac{1900-600}{300})="

"=P(-1<Z<1)=P(Z<1)-P(Z<-1)=0.6826."

Number of farms: "N=78*0.6826\approx53."


b) "P(X<1900)=P(Z<\frac{1900-1600}{300})=P(Z<1)=0.8413."